OpenStudy (anonymous):

[Properties of Negitive Exponents] - Simplify. Your answer should only contain positive exponents. 1) 4r^-3 2r^2 ---- 2) 4n^4 2n^-3

OpenStudy (anonymous):

1) rearrange, 4*2*r^-3r^2 = 8r^-1 = 8/r^1 = 8/r 2) 2*4*n^4n^-3 = 8*n^1= 8r

OpenStudy (anonymous):

Thank you! But what does the '*' mean :s?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

Thanks so much c:

OpenStudy (anonymous):

youre welcome